# ODEStream: A Buffer-Free Online Learning Framework with ODE-based Adaptor for Streaming Time Series Forecasting

### v1

@@ 4.2 Overview of ODEStream
Online learning is more complex than batch learning, where the model learns feature representations from sample data and applies them to unseen data. The primary objective of the streaming model is to forecast future values using specific horizontal windows over an extended period. Since the data is unavailable during training, the model does not have the opportunity to process and learn from the sequences over many iterations. In this case, when the data distribution eventually changes, and the sequences evolve over time, the learned features from former historical sequences may no longer be helpful in predicting future values. In the initial training phase of the model, often referred to as a warm-up, the model is pre-trained on historical data from the entire previous period before online learning commences. However, as discussed earlier, this data may have already changed or may change in the near future, so whatever the pre-trained model learns might not remain very relevant.
@@
For the ODEStream model warming-up phase, we do not warm up and train the model on a forecasting task with the goal of solely learning the relationship between historical data and future time steps. Instead, we follow a new approach. Since the prior data contains information about the distribution, dynamics, and temporal dependencies, our focus lies in initialising our model by utilising the available data to understand how this data evolves and how the dynamics change over time. To achieve this, we leverage the variational autoencoders (VAEs) (Liu et al., 2020) model as shown in the model framework in Figure 1. In contrast to traditional autoencoders, where an encoder layer $e$ maps the data $x$ to a latent vector $Z=e(x)$, in VAE, samples are transformed into an ideal data distribution $P(Z|X)$, and the encoder is used to generate the distribution rather than a single latent representation. For VAE, the model applies a prior constraint to the latent vector $Z$ using a multivariate Gaussian distribution $G$. The encoder $e$, parameterised with $\theta$, maps the data to a normal distribution rather than mapping it to a point in a latent space. The data $x$ distribution can be expressed as a multivariate Gaussian with mean $Z_{\mu}$ and log variance $Z_{\sigma}$ vectors where the probability function of this Gaussian distribution is:
$q(x|z)=\mathcal{N}(Z_{\mu},Z_{\sigma}),$ (3)
@@
here, $q(x|z)$ is the conditional distribution of the data $x$ given the latent variable $z$, and $\mathcal{N}$ is the Gaussian distribution with mean $Z_{\mu}$ and standard deviation $Z_{\sigma}$. Following that, the latent representation $z$ is sampled from the standard Gaussian distribution as:
$Z=z_{\mu}+\epsilon\bigodot\sqrt{exp(0.5\cdot z_{\sigma})}~{}~{},$ (4)
@@
where $\epsilon$ is a random sample drawn from a standard normal distribution $\mathcal{N}(0,1)$. The sampled latent variable $z$ is then fed to the decoder, which generates the final output $y$ by sampling the latent variable $z$ and mapping it to the final output $y$ as:
$p(x|z)=\mathcal{N}({z_{\mu}}^{\prime},{z_{\sigma}}^{\prime}),$ (5)
where ${z_{\mu}}^{\prime}$ and ${z_{\sigma}}^{\prime}$ are the mean and variance of the output data distribution.
@@
To be able to obtain the dynamics and learn how the sequence distribution and dynamics change over time, we use neural ODEs. As ODEs can learn the dynamics over time, they are suitable for tracking changes in time series data. The RNN encoder processes the input sequence $x$ to obtain the hidden state $h_{0}$. The hidden state is then mapped to the mean $Z_{\mu}$ and log variance $Z_{\sigma}$ discussed above using a linear layer. The latent variable $z$ is sampled from a Gaussian distribution as:$z_{t0}\sim\mathcal{N}({z_{\mu}},{z_{\sigma}})$
@@
Given a sequential data sequence $\{x_{1},x_{2},\cdots,x_{T}\}$ with a sequence length of $T$, a neural RNN encoder processes each time step $t$ to produce a hidden state $h_{t}$. The encoder estimates the parameters $Z_{\mu}$ and $Z_{\sigma}$ of the variational posterior distribution $p(z_{t}|x_{t})$ for the latent variable $z_{t}$. Subsequently, a fully connected layer maps the hidden state to $h_{0}$, and samples $z_{t0}$ using the following equation:
$z_{t0}\sim p(z_{t0}|x_{t0},\ldots,x_{T};t_{0},\ldots,t_{T};\theta_{p})=\mathcal{N}(z_{t0}|z_{\mu_{t}},z_{\sigma_{t}})$ (6)
@@ 4.4 Online Learning
After the initial training phase, where the neural ODEs model learns from historical data, the model switches into the online learning phase. In this phase, the trained model is deployed to serve as a dynamic adaptive network to provide continuous analysis for new streaming sequences as they emerge as in Figure 2. (For simplicity, we denote the sequence of $X=\{x_{i+1},x_{i+2},x_{i+3},\dots\}$, as $X=\{x_{1},x_{2},x_{3},\dots,x_{j},\dots\}$, where j represents the time point after $i$ assuming that $i$ is the last point of the data used for initial training.) The primary role of the network now is two-fold: first, to generate forecasts based on the recent available data $x_{j}$, and second, to adapt to changes in the data distribution. This adaptation is driven not only by prediction accuracy but also by the evolution of the data distribution, which will be discussed in Section 4.5. The model is meticulously designed to facilitate an ongoing understanding of how the underlying dynamics change over time. This is achieved by continuously incorporating new data sequences using the neural ODEs, allowing the model to learn and adapt to shifts in the data distribution. Particularly, the pre-trained neural ODE weights and architecture will be loaded, and the model will again used to model the dynamics of the latent space $z(j)$ of the streaming sample $x_{j}$where j is the last available value of the stream. Then, the latent space is linearly mapped to the hidden space $(h_{j})$. Finally, the hidden space is linearly mapped to the output space $(y_{j})$ as:
@@
Based on the learned dynamics, the model parameters $\theta$ are updated as $\theta_{\text{new}}=\theta_{\text{old}}-\eta\cdot\nabla_{\theta}\text{Loss}$
Figure 2: The continual learning process of ODEStream, The model processes the incoming observation using the pre-trained VAE-ODE and continuously changes its parameters based on new dynamics
Temporal Isolation Layer: To further enhance the model’s adaptability and responsiveness, we incorporated it with a specialised component known as the temporal isolation layer. This neural network serves two purposes: $(i)$ Focus on new information, as it resembles a dynamic “window” into the most recent data samples. Its primary objective is to extract and learn from the unique temporal patterns and features embedded within the recent data. By closely monitoring the look-back windows and forecast horizon of the data, this layer ensures that the model effectively captures emerging trends and patterns. $(ii)$ Avoid historical bias where, in addition to the focus on new information, we also manage the influence of past patterns learned by the pre-trained model that may have become obsolete or inaccurate. By independently processing and learning from the most recent data samples, this layer reduces the risk of the model being unduly guided by outdated historical data, ensuring its predictions remain relevant and reliable. The input of this layer is the current stream sequence with a specific look back window (lag) $(x_{j-lag},..x_{j})$, where j is the most recent time point of a stream. As shown in Equation 11 and 12, the temporal isolation layer $(f)$ generates the latent representation of that sample $h_{j}$, and a concatenate layer $(g)$ will combine the latent samples learned $h_{j}^{\prime}$ along with the neural ODE representation $h_{j}$ of that sample to generate the final prediction. $c_{j}$ is a cell state as we are using LSTM for this layer.
@@ 5 Experiments
  • What is the model adaptability to incoming data, how does it evolve over extended periods of streaming and, how does the concept drift in the data influence the performance?
@@ 5.1.2 Baseline
We address a set of baselines for continual learning time series forecasting. As FSNET (Pham et al., 2023) is the most recent state-of-the-art model, we consider it as our main baseline along with its variant naive that showed better performance for one step ahead forecasting. We also include a basic buffer-based model (Chaudhry et al., 2019b) Experience Replay (ER). ER enhances the OnlineTCN by incorporating an episodic memory that preserves past samples, subsequently interleaving them during the learning of more recent ones. Additionally, we consider a recent variant of ER, Dark experience replay DER++ (Buzzega et al., 2020), that Enhances the conventional ER by introducing a knowledge distillation loss on the previous logits.
5.1.3 Setup
For all the experiments, we applied non-shuffled splitting to divide the data into a training set (25% ) for initialising the model (warm-up training) and a streaming dataset (75%) for online testing and training. We rescaled (normalised) the features using a standard normal distribution that fits on the training set only. We used a look-back window of 24 for our model and the baselines, while the learning rate was set to 0.001, and we employed the Adam optimiser for model optimisation. For warm-up, we used a batch size of 64. To simulate a streaming environment for online learning, the batch size and epoch were set to one.
@@ 5.2 Evaluation Results
In the case of the ECL dataset, which is naturally univariate data, we conducted univariate forecasting by analysing the historical data of a specific variable (a single client). All other datasets are multivariate data, where we leverage multiple features to predict future values of a single target variable as multivariate forecasting tasks. For all datasets, we used all available features to predict a set of future values for these features for multivariate-multivariate forecasting. It is important to note that while reviewing the baseline models and the proposed model described in (Pham et al., 2023), we observed that the reported performance metrics were primarily for the multivariate-multivariate forecasting task where a multivariate predict multivariate. However, we extended our analysis to include the prediction performance for other tasks as well. Upon closer examination, the performance shown by the baseline models was not as high as anticipated in tasks other than multivariate-multivariate forecasting. For this task, the performance of our method closely approximated that of the baseline models. However, we achieved a substantial enhancement in the performance of univariate forecasting tasks when compared to the baseline models. Figure 3 and 4 illustrate the performance of ODEStream, FSNet and DER++ over a massive period of streaming, and we can see how well ODEStream and the baseline models adapt to gradual changes in the datasets through continuous learning. It is evident that FSNet and DER++ struggle to adapt to all the data changes over an extended period, resulting in predictions that diverge significantly from ground truth values for most of the dataset by the end of the streaming sequence. Conversely, ODEStream continuously adapts to new dynamics across the entire duration, which explains why it achieved lower cumulative MSE values. It is also worth mentioning that DER++ outperforms in multivariate forecasting due to the complex and non-stationary nature of the relationships between different variables. DER++ leverages a replay buffer that stores a small subset of past data points, allowing the model to rehearse past experiences. This mechanism helps maintain performance on previous data distributions while adapting to new patterns, which is crucial for handling the intricacies of multivariate time series data.
@@ 7.2 Computational Resources
In Figure 8, we show the required resources in terms of time and memory. In the first three plots, we present the required time needed to process a stream of sequences for each dataset. We have added the FSnet required time as a reference to show the improvement and provide insights into the efficiency of ODEStream. For example, Figure 8(a) shows the average time per second required to process one stream sequence. ODEStream requires 88% less on average. In Figure 8(d), the average memory usage in megabytes is calculated using the ‘psutil qlibrary for each process.

### v2

@@ 4.3 Obtaining Prior Knowledge (Model warm-up)
Online learning is more complex than batch learning, where the model learns feature representations from sample data and applies them to unseen data. The primary objective of the streaming model is to forecast future values using specific horizontal windows over an extended period. Since the data is unavailable during training, the model does not have the opportunity to process and learn from the sequences over many iterations. In this case, when the data distribution eventually changes, and the sequences evolve over time, the learned features from former historical sequences may no longer be helpful in predicting future values. In general, during the initial training phase of a continual learning model, often referred to as the warm-up phase, the model is pre-trained on historical data from the entire previous period before online learning commences. However, as discussed earlier, this data may have already changed or may change in the near future, so whatever the pre-trained model learns may not remain very relevant in the future. Therefore, relying solely on past data may not fully capture the evolving nature of the streaming sequences.
@@
For the ODEStream model warming-up phase, we do not warm up and train the model on a forecasting task with the goal of learning the relationship between historical data and future time steps. Instead, we follow a new approach. Since the prior data contains information about the distribution, dynamics, and temporal dependencies, our focus lies in initialising our model by utilising the available data to understand how this data evolves and how the dynamics change over time. To achieve this, we leverage the variational autoencoders (VAEs) (Liu et al., 2020) model as shown in the model framework in Figure 1. In contrast to traditional autoencoders, where an encoder layer $e$ maps the data $x$ to a latent vector $Z=e(x)$, in VAE, samples are encoded into a distribution $q(Z|X)$ over the latent variable $z$, rather than being transformed into a single latent representation. The main idea behind a VAE involves two key components; first, an encoder encodes the input data $x$ into a latent variable $z$, capturing its representation distribution through a probabilistic approach parametrized with $\theta$ (Kingma & Welling, 2014), next, a decoder reconstructs the original input data using the latent variable $z$.
In detail, the encoder $e$, which parametrised by $\theta$, produces the outputs: the mean $Z_{\mu}$ and log variance $Z_{\sigma}$ of a normal distribution, typically assuming a prior $p(z)$ as a standard multivariate Gaussian distribution, denoted as $\mathcal{N}(0,1)$. The following equation represents the probability function of the Gaussian distribution:
$q(z|x)=\mathcal{N}(Z_{\mu},Z_{\sigma}),$ (3)
@@
here, $q(z|x)$ is the conditional distribution of the data $x$ given the latent variable $z$, and $\mathcal{N}$ is the Gaussian distribution with mean $Z_{\mu}$ and standard deviation $Z_{\sigma}$. The latent variable $z$ is sampled from the reparameterized distribution as:
$Z=z_{\mu}+\epsilon\bigodot exp(0.5\cdot z_{\sigma}),$ (4)
@@
where $\epsilon$ is a random sample drawn from a standard normal distribution $\mathcal{N}(0,1)$. This sampled latent variable $z$ is then fed into the decoder, which generates the final output $y$ by expressing p(x|z) as:
$p(x|z)=\mathcal{N}({{\mu}_{y}},{{\sigma}_{y}^{2}}),$ (5)
where $\mu_{y}$ and $\sigma_{y}^{2}$ are the parameters of the decoder.
@@
To obtain the dynamics and learn how the sequence distribution changes over time, we use neural ODEs. As ODEs can learn the dynamics over time, they are suitable for tracking changes in time series data. The RNN encoder processes the input sequence $x$ to obtain the hidden state $h_{0}$. The hidden state is then mapped to the mean $Z_{\mu}$ and log variance $Z_{\sigma}$ discussed above using a linear layer. The latent variable $z$ is sampled from a Gaussian distribution as: $z_{t0}\sim\mathcal{N}({z_{\mu}},{z_{\sigma}})$
@@
Given a sequential data sequence $\{x_{1},x_{2},\cdots,x_{T}\}$ with a sequence length of $T$, a neural RNN encoder processes each time step $t$ to produce a hidden state $h_{t}$. The encoder estimates the parameters $Z_{\mu}$ and $Z_{\sigma}$ of the variational posterior distribution $q(z_{t}|x_{t})$ for the latent variable $z_{t}$. Subsequently, a fully connected layer maps the hidden state to $h_{0}$, and samples $z_{t0}$ using the following equation:
$z_{t0}\sim q(z_{t0}|x_{t0},\ldots,x_{T};t_{0},\ldots,t_{T};\theta_{p})=\mathcal{N}(z_{t0}|z_{\mu_{t}},z_{\sigma_{t}})$ (6)
@@ 4.4 Online Learning
Following the initial training phase, during which the neural ODE model learns from historical data, the model transitions into online learning. In this phase, the trained model is deployed as an adaptive dynamic network, enabling it to continuously process and analyse new streaming sequences as they become available (as illustrated in Figure 2). For simplicity, we denote the sequence $X=\{x_{i+1},x_{i+2},x_{i+3},\dots\}$, as $X=\{x_{1},x_{2},x_{3},\dots,x_{j},\dots\}$, where $j$ represents the time point following $i$, assuming $i$ denotes the last data point utilized in the initial training phase.
The main role of the adaptive network is twofold: first, to generate forecasts based on the recent available data $x_{j}$, and second, to adapt to changes in the data distribution. This adaptation is driven not only by prediction accuracy but also by the evolution of the data distribution, which will be discussed in Section 4.5.
The model architecture is intentionally designed to support a continual understanding of how the underlying data dynamics evolve over time. This achieved by continuously integrating new data sequences through the neural ODE framework, allowing the model to detect and adjust to shifts in data distribution. Specifically, the pre-trained neural ODE weights and structure are reloaded to model the latent space dynamics $z(j)$ of the streaming sample $x_{j}$, where $j$ denotes the most recent value in the data stream. The latent space representation $z(j)$ is then linearly mapped to a hidden space $(h_{j})$, which subsequently maps linearly to the output space $(y_{j})$ , as defined by:
@@
Based on the learned dynamics, the model parameters $\theta$ are updated iteratively as $\theta_{\text{new}}=\theta_{\text{old}}-\eta\cdot\nabla_{\theta}\text{Loss}$.
Figure 2: The continual learning process of ODEStream. The model processes the incoming observation using the pre-trained VAE-ODE and continuously changes its parameters based on new dynamics.
Temporal Isolation Layer: To further enhance the model’s adaptability and responsiveness, we incorporated it with a specialised component known as the temporal isolation layer. This neural network serves two purposes: $(i)$ Focus on new information, as it resembles a dynamic “window” into the most recent data samples. Its primary objective is to extract and learn from the unique temporal patterns and features embedded within the recent data. By closely monitoring the look-back windows and forecast horizon of the data, this layer ensures that the model effectively captures emerging trends and patterns. $(ii)$ Avoid historical bias, where, in addition to the focus on new information, we also manage the influence of past patterns learned by the pre-trained model that may have become obsolete or inaccurate. By independently processing and learning from the most recent data samples, this layer reduces the risk of the model being unduly guided by outdated historical data, ensuring its predictions remain relevant and reliable.
The input of this layer is the current stream sequence with a specific look back window (lag) $(x_{j-lag},\dots,x_{j})$, where $j$ is the most recent time point of a stream. As shown in Equation 11 and 12, the temporal isolation layer $(f)$ generates the latent representation of that sample $h_{j}^{\prime}$, and a concatenate layer $(g)$ combines the latent representation of the samples $h_{j}^{\prime}$ along with the neural ODE-based representation $h_{j}$ of that sample to generate the final prediction. $c_{j}$ is a cell state as we employ an LSTM for this layer
@@ 4.5 Model Regularisation
1 Input: Pre-trained NeuralODE model, streaming sequence $X=\{x_{1},x_{2},\dots,x_{j}\}$, previously learned parameters $\theta$
2for each time step $j$ in $X=\{x_{1},x_{2},\dots,x_{j}\}$ do:
3  $z(j)=\text{NeuralODE}(z_{0},j)$
4  $h_{j}=\text{Linear}(z(j))$
5  $h_{j}^{\prime},c_{j}\leftarrow f(x_{j},h_{j-1},c_{j-1})$
6  $h_{j}^{\prime\prime}=g(h_{j}^{\prime},h_{j})$
7  $y_{j+1}=\text{Linear}(h_{j}^{\prime\prime})$
8Update model parameters
9  $\mathcal{L}_{\text{KL}}=-0.5\sum_{i=1}^{N}\left(1+\log(Z_{\sigma_{i}}^{2})-Z_{\mu_{i}}^{2}-Z_{\sigma_{i}}^{2}\right)$
10  $\mathcal{L}_{\text{L1}}=0.01\sum_{i=1}^{N}|z_{i}|$
11  $\text{Loss}=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{L1}}$
12  $\theta_{\text{new}}=\theta_{\text{old}}-\eta\cdot\nabla_{\theta}\text{Loss}$
13end for
Output: Updated model parameters $\theta$ and final prediction $y_{j+1}$
Algorithm 1 ODEStream - Online Learning
Algorithm 1 outlines the ODEStream online learning algorithm, in summary, the algorithm employs a pre-trained Neural VAE-ODE model within an online learning framework, processing the streaming data sequence $X=\{x_{1},x_{2},\dots,x_{j}\}$ and updating the model parameters as new data becomes available. The ODEStream algorithm operates iteratively for each incoming data sample $x_{j}$, integrating both the neural ODE representation and the temporal isolation layer to generate predictions and update model parameter. Initially, the model calculates a latent representation $z(j)$ of the streaming sample using the NeuralODE function, followed by a linear mapping to hidden and output spaces $h_{j}$. Simultaneously, the temporal isolation layer generates an independent latent representation $h_{j}^{\prime}$ for the new sample, isolating recent patterns and minimizing reliance on potentially outdated historical data. At each time step, the algorithm combines the representations $h_{j}$ and $h_{j}^{\prime}$ using a concatenation layer to produce an updated hidden state $h_{j}^{\prime\prime}$, yielding the final prediction $y_{j+1}$. The model parameters $\theta$ are then updated via gradient descent, optimizing with a composite loss function incorporating mean squared error (MSE), KL-divergence $\mathcal{L}_{\text{KL}}$, and L1 regularization $\mathcal{L}_{\text{L1}}$. This procedure ensures that the model adapts to new data patterns while preserving relevant insights from previous training.
@@ 5 Experiments
  • How adaptable is the model to incoming data? How does it evolve over extended periods of streaming data, and how does concept drift in the data influence its performance?
@@ 5.1.2 Baseline
In addition to a non-continual learning baseline using a basic RNN layer, we evaluated a set of baselines for continual learning time series forecasting. As FSNET (Pham et al., 2023) is the most recent state-of-the-art model, we consider it as our main baseline along with its variant naive that showed better performance for one step ahead forecasting. We also include a basic buffer-based model (Chaudhry et al., 2019b); Experience Replay (ER). ER enhances the OnlineTCN by incorporating an episodic memory that preserves past samples, subsequently interleaving them during the learning of more recent ones. Additionally, we consider a recent variant of ER, Dark experience replay DER++ (Buzzega et al., 2020), that Enhances the conventional ER by introducing a knowledge distillation loss on the previous logits.
5.1.3 Setup and Experimental details
In the encoder part of the VAE, we utilize RNN model consisting of one GRU layer with a hidden dimension of 64, followed by a linear layer that outputs a latent dimension of 64. The decoder comprises three main components: a Neural-ODE that initializes the latent state $z$, a linear layer that maps the latent state to the hidden dimension of 64, and a second linear layer that maps the hidden representation to the final output. The dynamics of the system are integrated over time using the Euler method, with the step size determined based on the time duration from the initial time $t_{0}$ to the final time $t_{1}$, constrained by a maximum step size of 0.05.
For all the experiments, we applied non-shuffled splitting to divide the data into a training set (25% ) for initialising the model (warm-up training) and a streaming dataset (75%) for online testing and training. We rescaled (normalised) the features using a standard normal distribution that fits on the training set only. We used a look-back window of 24 for our model and the baselines, while the learning rate was set to 0.001, and we employed the Adam optimiser for model optimisation. During the warm-up phase, we utilized a batch size of 64 and implemented early stopping after 10 epochs. To simulate a streaming environment for online learning, the batch size and epoch were set to one.
@@ 5.2 Evaluation Results
In the case of the ECL dataset, which is naturally univariate data, we conducted univariate forecasting by analysing the historical data of a specific variable (a single client). All other datasets are multivariate data, where we leverage multiple features to predict future values of a single target variable as multivariate forecasting tasks. For all datasets, we used all available features to predict a set of future values for these features for multivariate-multivariate forecasting. It is important to note that while reviewing the baseline models and the proposed model described in (Pham et al., 2023), we observed that the reported performance metrics were primarily for the multivariate-multivariate forecasting task where a multivariate predict multivariate. However, we extended our analysis to include the prediction performance for other tasks as well. Regarding the non-continual learning baseline (RNN model) the performance is suboptimal due to the fact that these models are not designed for continual learning and are unsuitable for handling streaming environments effectively. Upon closer examination, the performance shown by the other baseline models was not as high as anticipated in tasks other than multivariate-multivariate forecasting. For this task, the performance of our method closely approximated that of the baseline models. However, we achieved a substantial enhancement in the performance of univariate forecasting tasks when compared to the baseline models. Figure 3 and 4 illustrate the performance of ODEStream, FSNet and DER++ over a massive period of streaming, and we can see how well ODEStream and the baseline models adapt to gradual changes in the datasets through continuous learning. It is evident that FSNet and DER++ struggle to adapt to all the data changes over an extended period, resulting in predictions that diverge significantly from ground truth values for most of the dataset by the end of the streaming sequence. Conversely, ODEStream continuously adapts to new dynamics across the entire duration, which explains why it achieved lower cumulative MSE values. It is also worth mentioning that DER++ outperforms in multivariate forecasting due to the complex and non-stationary nature of the relationships between different variables. DER++ leverages a replay buffer that stores a small subset of past data points, allowing the model to rehearse past experiences. This mechanism helps maintain performance on previous data distributions while adapting to new patterns, which is crucial for handling the intricacies of multivariate time series data.
In addition to the main experiment, Table 2 presents the results of the multi-horizon predictions for 7 and 24 steps across five datasets. The performance remains stable over longer prediction horizons, with no significant degradation in accuracy. Notably, even at 7 and 24 steps, the model consistently outperforms the baseline models from previous experiments.
Table 2: Multi-horizon prediction results.
steps ECL ETTH1 ETTH2 ETTM1 WTH
1 0.1173 0.0594 0.164 0.0625 0.0441
7 0.403 0.1006 0.4009 0.5997 0.1089
24 1.003 0.1909 0.3751 0.09186 0.133
@@ 7.2 Computational Resources
In Figure 8, we show the required resources in terms of time and memory. In the first three plots, we present the required time needed to process a stream of sequences for each dataset. We have added the FSnet required time as a reference to show the improvement and provide insights into the efficiency of ODEStream. For example, Figure 8(a) shows the average time per second required to process one stream sequence. ODEStream requires 88% less time on average compared with the baseline model. In Figure 8(d), the average memory usage in megabytes is calculated using the psutil library for each process.